OpenStudy (anonymous):

Suppose that a fourth order differential equation has a solution y=3e^(4x)xcos(x). (a) Find such a differential equation, assuming it is homogeneous and has constant coefficients.

OpenStudy (primeralph):

differentiate this up to 4 times and make some expression with it.

OpenStudy (anonymous):

That's one way of doing it.

OpenStudy (anonymous):

still confusing!

OpenStudy (anonymous):

So you have to find a 4th order ODE that looks like this: \[ay^{(4)}+by'''+cy''+dy'+ey=0\] The characteristic equation is then \[ar^4+br^3+cr^2+dr+e=0\] You know one solution to the ODE is \(y=3e^{-4x}x\cos x\). This gives you a few bits of information: ∙ \(e^{\lambda x}\cos \mu x~\Rightarrow~r=\lambda+\mu i\) is a solution to the characteristic equation. (Since complex roots come in conjugate pairs, you also know that \(r=\lambda-\mu i\) is also a solution to the characteristic equation.) ∙ The power of \(x\) in the solution indicates a repeated root. So, you have to find \(a,b,c,d,e\) such that \[ar^4+br^3+cr^2+dr+e=0\] has roots \(r=-4+i\), with multiplicity 2.